OpenStudy (anonymous):

The radius of a sphere , initially zero, increases at the rate of 6 ft. per sec. Find how fast the volume increases after 1/4 sec.. pls help

OpenStudy (anonymous):

\[Volume=\frac{ 4 }{ 3}\pi r^{3}\] Initially the volume is 0. After 1/4 second, the radius will be 6/4 = 1.5 feet Use the formula and find the volume which should be \[\frac{ 9 }{ 2 }\pi\]